OpenStudy (anonymous):

9(m+2)=-6(m+7) Show Work.

OpenStudy (shaik0124):

9m+18=-6m-42 9m+18-18=-6m-42-18 9m=-6m-60 9m+6m=-6m+6m-60 15m=-60 m=-60/15 m=-4

OpenStudy (shaik0124):

did u understood@xG...

OpenStudy (anonymous):

no because there isn't anything in parentheses.

OpenStudy (shaik0124):

no solve in the way i have done it simple to understand too